# Static spheres and Aschenbach effect for black holes in massive gravity

### v1

@@ Abstract
In this paper, we study the trajectories of massive and massless particles in four dimensional static and spherically symmetric black holes in dRGT massive gravity theory via phase-plane analysis and point out several novel features. In particular, we show the existence of a static sphere, a finite radial distance outside the black holes in these theories, where a massive particle can be at rest, as seen by an asymptotic zero angular momentum observer. Topological arguments show that the stable and unstable static spheres, which come in pairs, have opposite charges. In the presence of angular momentum, we first study the behaviour of massless particles and find the presence of stable and unstable photon spheres in both neutral and charged black holes. Subsequently, we study the motion of massive test particles around these black holes, and find one pair of stable and unstable time-like circular orbits (TCOs), such that the stable and unstable TCO’s are disconnected in certain regions. Computing the angular velocity $\Omega_{\text{\tiny CO}}$ of the TCOs, measured by a static observer at rest, shows the unusual nature of its monotonic increase with the radius of TCO, near the location of stable photon sphere. This confirms the existence of Aschenbach effect for spherically symmetric black holes in massive gravity, which was only found to exist in rapidly spinning black holes, with the only other exception being the rare example of gravity coupled to quasi-topological electromagnetism.
@@ 1 Introduction
Investigation of orbits of particles is ubiquitous for uncovering novel observable gravitational phenomena and testing various features of compact objects in Einstein gravity as well as modified theories of gravity. Applications range from situations involving strong and weak gravitational effects around black holes, such as ring down in a binary system of black holes[1], shadows[2, 3], and to other interesting phenomena in four and higher dimensions, with potential observational signatures [4, 5, 6, 7, 8, 9, 10, 11]. Stable circular orbits of massive particles around black holes can be used to study accretion discs, whereas unstable ones provide information about shadows. Certain physical quantities, such as Energy, angular momentum and angular velocity are generally examined to extract information about the orbital dynamics [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Another related development which could be worthwhile to explore, especially, for testing modified theories of gravity, concerns an intriguing effect found by Aschenback[33], coming from the study of radial gradient of orbital velocity around rotating black hole space-times. Let us elaborate. Generally, in Newtonian mechanics, or even while studying orbits of massive particles around spherically symmetric and/or rotating black holes, the angular velocity increases monotonically with decreasing radius. This behaviour also continues to hold for small values of rotation $a$ and large orbit radius $r$. What was found in[33] is that, for rapidly spinning black holes with $a>0.9953$ (but still below the Thorne limit[34] $a=0.998$), the radial gradient of orbital velocity can decrease with decreasing $r$ in a narrow window of orbital radius, before resuming its normal increasing behaviour. This effect is very specific to zero angular momentum observers (ZAMOs) or so called Bardeen observers, and the results appear to be commensurate with the high frequency quasi-periodic oscillations originating in the accretion disk of compact objects, such as from X-ray binaries. Considering the possibility of experimental observations, the Aschenback effect has received increased attention with investigations extended to more general Kerr black holes[35, 36, 37], as well as when the particles are charged[38], and/or rapidly spinning [39, 40].
@@
Now, since the Aschenbach effect occurs for test particles in orbits around rapidly spinning black holes[33, 35], as observed by ZAMOs, it was not expected to be present when the black hole is non-rotating (or slowly rotating). At least, this is what can be deduced for spherically symmetric black holes in Einstein gravity. However, recently, in the aforementioned model of Einstein gravity coupled to nonlinear electrodynamics, the Aschenback effect was found to be present [29], aided by the occurrence of static spheres in a certain narrow range of charge to mass ratio of the black hole. As emphasised earlier, due to the possibility of potential experimentally observational signatures, the Aschenback effect should be explored further, in more general settings, which could give rise to clues to the possible nature of the compact object. In this paper, we find a novel example where the Aschenback effect can be present in a theory of massive gravity, which contains static and spherically symmetric black holes, with a rich structure of stable/unstable circular orbits of test particles. Our study is also expected to be helpful while exploring various lower bounds on physical quantities related to orbits of black holes[41], particularly, in massive gravity.
@@
Following are some broad motivations for considering a massive theory of gravity. Einstein’s gravity has had high success with experimental observation of several of its predictions, buoyed further with the most recent evidences coming from the LIGO collaboration [42, 43] pertaining to gravitational wave signals. Notwithstanding these triumphs, there continue to be other phenomena of paramount importance, foremost being the accelerated expansion of the universe, and the longstanding cosmological constant problem, among others, necessitating exploration of modified theories of gravity. One of the possibilities explored concerns massive graviton theories, looked into with motivation from hierarchy problems and strong connections with quantum theory of gravity [44, 45], which is within the ambit of recent observations [46], where novel limits on the graviton mass are obtained. We should mention that, massive theories of gravity have a good history, with the first models being constructed as early as 1939 by Firez and Paullo [47]. These models of course have now undergone considerable modifications, most notably the the advent of new massive gravites[48, 49, 50, 51] which are being investigated actively[52, 53, 54, 55, 56, 57, 58]. Black hole solutions in these theories have interesting thermodynamic phases [59, 60, 61, 62] with wide ranging cosmology/astrophysics applications, which should be helpful in identifying deviations from Einstein’s theory [63, 64, 65, 66, 67, 68, 69]. Of interest to this work is an interesting class of massive theory of gravity proposed in [70], which has a well formulated holographic dual description. The model in [70] has a singular metric, with other useful features of stability and absence of ghosts[71], together with the presence of black holes[72, 73, 74, 75, 76]. There are considerable advantages in constructing such massive theories of gravity, which have the necessary ingredients for addressing some of the aforementioned issues as elaborated in [77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89]: to mention one, it may be possible to understand the current observations emanating from dark matter [90] and also correlating with the accelerating expansion of our universe where the requirement of dark energy component could be relaxed [91, 92]. Let us also mention in passing, the continued efforts in embedding massive theories of gravity in the string theory framework [93]. Some recent developments in general theories of gravity are also on the thermodynamic front, where the cosmological constant is taken to a variable, giving rise to the concept of pressure and black hole chemistry program (see[94] and references therein), which has been explored extensively in massive gravity theories [95, 76, 75, 96, 97, 98, 99, 100, 101, 102]. The issue of black hole microstructures in these theories have also been explored from the point of view of thermodynamic geometry with interesting results [103, 104, 105, 106, 107, 108, 109].
@@
Rest of the paper is organised as follows. In section-(2), we present a brief overview of the static and spherically symmetric black holes in massive gravity in four dimensions. These black holes, often endowed with multiple horizons, have rich geometric structures. Consequently, motions of massive and massless particles around these black holes possess diverse trajectories, including that of the static spheres. To explore the nature of these trajectories, in section-(3), we develop a phase-plane analysis of the system. We find the fixed points and complete phase portraits, dictated by those fixed points. Owing to the fact that the black holes in massive gravity have several free parameters, our study in this section is not exhaustive. We fix some of these parameters and vary the rest in a way that captures qualitatively different geodesics. Subsequent section is devoted to examining one class of these geodesics associated with the massive particles, namely the static spheres. In section-(4.1), we present the conditions for the existence of circular orbits for massive test particles when the orbital angular momentum vanishes, and then proceed to count the number of such static spheres, based on topological arguments. Exploiting the presence of static spheres and stable photon spheres in the massive gravity, in section-(4.3), we study the Aschenback effect, where the angular velocity of a time-like circular orbit is found to be increasing with its radius coordinate under certain conditions. Finally, we conclude in section-(5).
@@ 3.1 Fixed points and phase portraits for massless particle
$A:(z,x)=(0,\frac{1}{6M}(1+\zeta-\sqrt{1+6M\gamma+2\zeta+\zeta^{2}}),~{}~{}{\rm and}$
$B:(z,x)=(0,\frac{1}{6M}(1+\zeta+\sqrt{1+6M\gamma+2\zeta+\zeta^{2}}),$ (3.10)
@@
Figure 2: The figures show the phase plane portraits of massless particle outside an uncharged black hole. We vary $(\alpha,\beta)$ and keep $M=h=m_{g}=1,Q=0$. Top left: the fixed point outside the horizon is a center producing closed orbits. We have chosen $(\alpha,\beta)=(-4,2.9)$. This point lies in the black-dotted region in Fig. (1). Top right: the fixed point now is a saddle which lies in the blue-dotted region of Fig.(1). We have taken $(\alpha,\beta)=(2,-2)$. Bottom row: these are for $(\alpha,\beta)=(2,4.1)$ (left) and $(8,3.1)$ (right), chosen from the red-dotted region of Fig. (1). In this region, both the fixed points are outside with the saddle being nearer to the horizon. In all these figures, the vertical lines drawn at extreme right represent the location of the outer horizon.
@@
[…]
@@ 3.2 Fixed points and phase portraits for massive particle
Figure 6: These figures show representative phase plane portrait of massive particle outside a charged black hole for $M=1,Q=.2,L=.5,m_{g}=h=1$. Left: A pair of nodes appear outside the outer horizon in the blue region – one is a saddle and the other being a center. Right: Typical portrait for $\alpha=8,\beta=4.1$. Nodes are marked with black dots. The vertical line on the right is the location of the outer horizon.
@@ 4.1 Existence of static spheres
[…]
@@ 5 Conclusions
Owing to the structure of the black holes in the massive gravity, the particle trajectories around these black holes could be diverse in nature. Therefore, via a phase-plane analysis, we first examined the fixed points and the phase portraits of these particles. We found that the fixed points, that lie outside the outer horizon, are either centers or saddle-nodes or both. Consequently, the trajectories offer closed orbits, including the circular one. Subsequently, we focussed on the possible existence of circular orbits for a massive test particle when its orbital angular momentum vanishes. These type of circular orbits are called as static spheres, where one may realise the construction of a static Dyson-like sphere [29]. We found that a pair of stable and unstable static spheres exists in both neutral and charged black hole backgrounds. Making use of a topological argument for static spheres, we showed that, in a general static and spherically symmetric black hole (charged or neutral) background in massive gravity, the stable and unstable static spheres always come in pairs, if they exist. The stable and unstable static sphere possess the topological charge $+1$ and $-1$, respectively. These topological classifications are found to be in agreement with the analysis done for a dyonic black hole in quasitopological electromagnetism [29]. Next, we searched for the existence of stable photon spheres. We observed that, for a given angular momentum $L$, there exists a pair of stable and unstable photon spheres in both neutral and charged black hole backgrounds. In [113], the presence of static spheres and stable photon spheres around a static and spherically symmetric black hole, signalled the occurrence of Aschenbach effect. Motivated by this, we focused on the properties of time-like circular orbits (TCOs) in order to see such effect in the massive gravity theory, with results distinct from Einstein gravity. In the case of dyonic black hole in quasi-topological electromagnetism, there exist two pairs of TCOs for a given angular momentum [113]. In contrast, for the neutral and charged black hole backgrounds in massive gravity considered here, we found only one pair of stable and unstable TCOs, such that the stable one is the outer TCO, and the unstable one is the inner TCO. Moreover, it is observed that certain TCOs have imaginary angular momentum, due to which the stable and unstable TCO regions are disconnected. We then proceeded by computing the angular velocity $\Omega_{\text{\tiny CO}}$ of the TCOs, measured by a static observer at rest. We found the monotonically increasing behaviour of $\Omega_{\text{\tiny CO}}$ with radius of the TCO, in the stable region of TCOs region, near the stable photon sphere. This then confirms the existence of Aschenbach effect for time-like particle motion around both neutral and charged black hole backgrounds in massive gravity. Together with the dyonic black hole in quasi-topological electromagnetism [113, 29], the neutral and charged black hole backgrounds in massive gravity studied in this work are the other examples for existence of static spheres and the Aschenbach effect as well. It would be interesting to find other systems, that show these properties. Finding a general condition, possibly using the topological arguments, which could allow prediction of static spheres and Aschenbach effect, would be nice to explore. Checking the occurrence of Aschenbach effect with a charged test particle for our system and also for dyonic black hole in quasi-topological electromagnetism [113], in addition to studying the effect of charge of the black hole, are possible avenues for future work. Investigating static spheres and Aschenbach effect from AdS/CFT point of view, possibly using some recently developed techniques [115, 116], would likely result in novel holographic understanding. We leave these issues for feature work.

### v2

@@ Abstract
In this paper, we study the trajectories of massive and massless particles in four dimensional static and spherically symmetric black holes in de Rham-Gabadadze-Tolley (dRGT) massive gravity theory via phase-plane analysis and point out several novel features. In particular, we show the existence of a static sphere, a finite radial distance outside the black holes in these theories, where a massive particle can be at rest, as seen by an asymptotic zero angular momentum observer. Topological arguments show that the stable and unstable static spheres, which come in pairs, have opposite charges. In the presence of angular momentum, we first study the behaviour of massless particles and find the presence of stable and unstable photon spheres in both neutral and charged black holes. Subsequently, we study the motion of massive test particles around these black holes, and find one pair of stable and unstable time-like circular orbits (TCOs), such that the stable and unstable TCO’s are disconnected in certain regions. Computing the angular velocity $\Omega_{\text{\tiny CO}}$ of the TCOs, measured by a static observer at rest, shows the unusual nature of its monotonic increase with the radius of TCO, near the location of stable photon sphere. This confirms the existence of Aschenbach effect for spherically symmetric black holes in massive gravity, which was only found to exist in rapidly spinning black holes, with the only other exception being the rare example of gravity coupled to quasi-topological electromagnetism.
@@ 1 Introduction
Investigation of orbits of particles is ubiquitous for uncovering novel observable gravitational phenomena and testing various features of compact objects in Einstein gravity as well as modified theories of gravity. Applications range from situations involving strong and weak gravitational effects around black holes, such as ring down in a binary system of black holes[1], shadows[2, 3], and to other interesting phenomena in four and higher dimensions, with potential observational signatures [4, 5, 6, 7, 8, 9, 10, 11]. Stable circular orbits of massive particles around black holes can be used to study accretion discs, whereas unstable ones provide information about shadows. Certain physical quantities, such as Energy, angular momentum and angular velocity are generally examined to extract information about the orbital dynamics [12, 13, 14, 13, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@
Another related development which could be worthwhile to explore, especially, for testing modified theories of gravity, concerns an intriguing effect found by Aschenbach[36], coming from the study of radial gradient of orbital velocity around rotating black hole space-times. Let us elaborate. Generally, in Newtonian mechanics, or even while studying orbits of massive particles around spherically symmetric and/or rotating black holes, the angular velocity increases monotonically with decreasing radius. This behaviour also continues to hold for small values of rotation $a$ and large orbit radius $r$. What was found in[36] is that, for rapidly spinning black holes with $a>0.9953$ (but still below the Thorne limit[37] $a=0.998$), the radial gradient of orbital velocity can decrease with decreasing $r$ in a narrow window of orbital radius, before resuming its normal increasing behaviour. This effect is very specific to zero angular momentum observers (ZAMOs) or so called Bardeen observers, and the results appear to be commensurate with the high frequency quasi-periodic oscillations originating in the accretion disk of compact objects, such as from X-ray binaries. Considering the possibility of experimental observations, the Aschenbach effect has received increased attention with investigations extended to more general Kerr black holes[38, 39, 40], as well as when the particles are charged[41], and/or rapidly spinning [42, 43].
@@
Now, since the Aschenbach effect occurs for test particles in orbits around rapidly spinning black holes[36, 38], as observed by ZAMOs, it was not expected to be present when the black hole is non-rotating (or slowly rotating). At least, this is what can be deduced for spherically symmetric black holes in Einstein gravity. However, recently, in the aforementioned model of Einstein gravity coupled to nonlinear electrodynamics, the Aschenbach effect was found to be present [32], aided by the occurrence of static spheres in a certain narrow range of charge to mass ratio of the black hole. As emphasised earlier, due to the possibility of potential experimentally observational signatures, the Aschenbach effect should be explored further, in more general settings, which could give rise to clues to the possible nature of the compact object. In this paper, we find a novel example where the Aschenbach effect can be present in a theory of massive gravity, which contains static and spherically symmetric black holes, with a rich structure of stable/unstable circular orbits of test particles. Our study is also expected to be helpful while exploring various lower bounds on physical quantities related to orbits of black holes[44], particularly, in massive gravity.
@@
Following are some broad motivations for considering a massive theory of gravity. Einstein’s gravity has had high success with experimental observation of several of its predictions, buoyed further with the most recent evidences coming from the LIGO collaboration [45, 46] pertaining to gravitational wave signals. Notwithstanding these triumphs, there continue to be other phenomena of paramount importance, foremost being the accelerated expansion of the universe, and the longstanding cosmological constant problem, among others, necessitating exploration of modified theories of gravity. One of the possibilities explored concerns massive graviton theories, looked into with motivation from hierarchy problems and strong connections with quantum theory of gravity [47, 48], which is within the ambit of recent observations [49], where novel limits on the graviton mass are obtained. We should mention that, massive theories of gravity have a good history, with the first models being constructed as early as 1939 by Firez and Pauli [50]. These models of course have now undergone considerable modifications, most notably the advent of new massive gravites[51, 52, 53, 54] which are being investigated actively[55, 56, 57, 58, 59, 60, 61]. Black hole solutions in these theories have interesting thermodynamic phases [62, 63, 64, 65] with wide ranging cosmology/astrophysics applications, which should be helpful in identifying deviations from Einstein’s theory [66, 67, 68, 69, 70, 71, 72] . There are considerable advantages in constructing massive theories of gravity [73] (see also the reviews [74, 75]), which have the necessary ingredients for addressing some of the aforementioned issues as elaborated in [76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88]: to mention one, it may be possible to understand the current observations emanating from dark matter [89] and also correlating with the accelerating expansion of our universe where the requirement of dark energy component could be relaxed [90, 91]. Let us also mention in passing, the continued efforts in embedding massive theories of gravity in the string theory and holographic framework [92, 93, 94]. Some recent developments in general theories of gravity are also on the thermodynamic front, where the cosmological constant is taken to a variable, giving rise to the concept of pressure and black hole chemistry program (see[95] and references therein), which has been explored extensively in massive gravity theories [96, 97, 98, 99, 100, 101, 102, 103, 104, 105]. The issue of black hole microstructures in these theories have also been explored from the point of view of thermodynamic geometry with interesting results [106, 107, 108, 109, 110, 111, 112].
With regard to the massive gravity theories considered in this work, we focus on the black hole solutions in the dRGT massive gravity theory in a four dimensional Einstein-Maxwell gravity coupled to nonlinear interaction terms giving the graviton a mass $m_{g}$ [113, 114, 115]. This theory includes a non dynamical reference metric $f_{\mu\nu}$, which breaks diffeomorphism invariance and may be singular [73]. However, the theory has several useful features of stability and absence of ghosts[116], together with the presence of black holes[117, 118, 119, 98, 97]. In particular, this set up is well suited for holographic discussions of momentum dissipation, as the framework in [73] describes a class of strongly coupled quantum field theories with broken translational symmetry, due to the Lorentz breaking graviton mass term. The later aspect is helpful, as due to this, the mass of the graviton can play a role similar to the one played by lattice in the holographic models of conductors. The conductivity in general possesses a Drude peak which in the massless gravity limit tends to a delta function. Thus, non-linear massive gravity theories require an auxiliary reference metric (for giving mass to gravitons) to yield a Boulware-Deser ghost-free theory [46, 74, 61]. In principle, a special theory of massive gravity for each choice of reference metric can be constructed, which is non-degenerate [61]. In the context of holographic duality, massive gravity theories in AdS space with a singular (degenerate) reference metric are helpful to model a class of strongly interacting quantum field theories having broken translational symmetry (that is, momentum dissipation). For instance, one can build a holographic model for normal conductors (with momentum dissipation) and finite DC conductivity [73, 120], in contrast with massless gravity theories, such as, Einstein and with higher derivatives, with infinite DC conductivity  [121, 122, 123, 124]. Furthermore, massive gravity theories with the feature of a singular reference metric can effectively capture different phases of known condensed matter systems having broken translational symmetry, such as solids, liquids and (perfect) fluids [99, 125, 126], with interesting applications to cosmological [127] and other holographic situations [73, 120, 128, 129, 130]. Of course, for a more viable description of massive gravity, the reference metric needs to be invertible and dynamical, which is the program pursued in bimetric gravity theories [89, 90, 91, 109, 110] (see also [131, 132, 133]). We should mention that, there are alternatives to using a reference metric $f_{\mu\nu}$, namely, the Stückelberg fields [54], where the holographic conductivity can be studied by regarding the fields as dynamical[99]. In the Stückelberg formalism of massive gravity, the diffeomorphism invariance seems to be restored. The methods and techniques developed in this work though are very general, and can be carried over to study these interesting massive gravity theories.
@@
Rest of the paper is organised as follows. In section-(2), we present a brief overview of the static and spherically symmetric black holes in massive gravity in four dimensions. These black holes, often endowed with multiple horizons, have rich geometric structures. Consequently, motions of massive and massless particles around these black holes possess diverse trajectories, including that of the static spheres. To explore the nature of these trajectories, in section-(3), we develop a phase-plane analysis of the system. We find the fixed points and complete phase portraits, dictated by those fixed points. Owing to the fact that the black holes in massive gravity have several free parameters, our study in this section is not exhaustive. We fix some of these parameters and vary the rest in a way that captures qualitatively different geodesics. Subsequent section is devoted to examining one class of these geodesics associated with the massive particles, namely the static spheres. In section-(4.1), we present the conditions for the existence of circular orbits for massive test particles when the orbital angular momentum vanishes, and then proceed to count the number of such static spheres, based on topological arguments. Exploiting the presence of static spheres and stable photon spheres in the massive gravity, in section-(4.3), we study the Aschenbach effect, where the angular velocity of a time-like circular orbit is found to be increasing with its radius coordinate under certain conditions. Finally, we conclude in section-(5).
@@ 3.1 Fixed points and phase portraits for massless particle
$A:(z,x)=(0,\frac{1}{6M}(1+\zeta-\sqrt{1+6M\gamma+2\zeta+\zeta^{2}})),~{}~{}{\rm and}$
$B:(z,x)=(0,\frac{1}{6M}(1+\zeta+\sqrt{1+6M\gamma+2\zeta+\zeta^{2}})),$ (3.10)
@@
Figure 2: The figures show the phase plane portraits of massless particle outside an uncharged black hole. We vary $(\alpha,\beta)$ and keep $M=h=m_{g}=1,Q=0$. Top left: the fixed point outside the horizon is a center producing closed orbits. We have chosen $(\alpha,\beta)=(-4,2.9)$. This point lies in the black-dotted region in Fig. (1). Top right: the fixed point now is a saddle which lies in the blue-dotted region of Fig.(1). We have taken $(\alpha,\beta)=(2,-2)$. Bottom row: the plots are for $(\alpha,\beta)=(2,4.1)$ (left) and $(8,3.1)$ (right), chosen from the red-dotted region of Fig. (1). In this region, both the fixed points are outside with the saddle being nearer to the horizon. In all these figures, the vertical lines drawn at extreme right represent the location of the outer horizon.
@@
Figure 3: Number of positive roots of the right equation of (3.17) are shown in the $(\alpha,\beta)$ plane using three distinct colours. Further, the regions with black dots $(P~{}{\rm and}~{}Q)$ hold only one of the fixed points outside the outer horizon. While in region $P$, the node is a center, a saddle appears at $Q$. Two fixed points are located outside the outer horizon in the other two dotted regions $(R~{}{\rm and}~{}S)$. In $R$, the saddle node is closer to the horizon and in $S$, the center. In the rest of the parameter space, either no fixed point arises outside the outer horizon or they are complex. We have set $M=h=m_{g}=1$, and $Q=0.2$.
@@ 3.2 Fixed points and phase portraits for massive particle
Figure 6: These figures show representative phase plane portrait of massive particle outside a charged black hole for $M=1,Q=0.2,L=.5,m_{g}=h=1$. Left: A pair of nodes appear outside the outer horizon in the blue region – one is a saddle and the other being a center. Right: Typical portrait for $\alpha=8,\beta=4.1$. Nodes are marked with black dots. The vertical line on the right is the location of the outer horizon.
@@ 4.1 Existence of static spheres
Figure 10: Static spheres around a charged black hole. We have set $M=m_{g}=h=1$ and $Q=0.2$. Left: Regions in blue show the values of $\alpha,\beta$ for which static spheres occur. Ranges of $\alpha,\beta$ are chosen only for the representative purpose. Middle: Fixed points, denoted in black, and the vector fields around those are shown. The chosen values of $(\alpha,\beta)$ are $(7,5)$. Shape of the effective potential following from (4.2) for $\alpha=7$ and $\beta=5$. The energy associated with the particle is $E=1.98$.
@@ 5 Conclusions
Owing to the structure of the black holes in the massive gravity, the particle trajectories around these black holes could be diverse in nature. Therefore, via a phase-plane analysis, we first examined the fixed points and the phase portraits of these particles. We found that the fixed points, that lie outside the outer horizon, are either centers or saddle-nodes or both. Consequently, the trajectories offer closed orbits, including the circular one. Subsequently, we focussed on the possible existence of circular orbits for a massive test particle when its orbital angular momentum vanishes. These type of circular orbits are called as static spheres, where one may realise the construction of a static Dyson-like sphere [32]. We found that a pair of stable and unstable static spheres exists in both neutral and charged black hole backgrounds. Making use of a topological argument for static spheres, we showed that, in a general static and spherically symmetric black hole (charged or neutral) background in massive gravity, the stable and unstable static spheres always come in pairs, if they exist. The stable and unstable static sphere possess the topological charge $+1$ and $-1$, respectively. These topological classifications are found to be in agreement with the analysis done for a dyonic black hole in quasitopological electromagnetism [32]. Next, we searched for the existence of stable photon spheres. We observed that, for a given angular momentum $L$, there exists a pair of stable and unstable photon spheres in both neutral and charged black hole backgrounds. In [134], the presence of static spheres and stable photon spheres around a static and spherically symmetric black hole, signalled the occurrence of Aschenbach effect. Motivated by this, we focused on the properties of time-like circular orbits (TCOs) in order to see such effect in the massive gravity theory, with results distinct from Einstein gravity. In the case of dyonic black hole in quasi-topological electromagnetism, there exist two pairs of TCOs for a given angular momentum [134]. In contrast, for the neutral and charged black hole backgrounds in massive gravity considered here, we found only one pair of stable and unstable TCOs, such that the stable one is the outer TCO, and the unstable one is the inner TCO. Moreover, it is observed that certain TCOs have imaginary angular momentum, due to which the stable and unstable TCO regions are disconnected. We then proceeded by computing the angular velocity $\Omega_{\text{\tiny CO}}$ of the TCOs, measured by a static observer at rest. We found the monotonically increasing behaviour of $\Omega_{\text{\tiny CO}}$ with radius of the TCO, in the stable region of TCOs region, near the stable photon sphere. This then confirms the existence of Aschenbach effect for time-like particle motion around both neutral and charged black hole backgrounds in massive gravity. Together with the dyonic black hole in quasi-topological electromagnetism [134, 32], the neutral and charged black hole backgrounds in massive gravity studied in this work are the other examples for existence of static spheres and the Aschenbach effect as well. It would be interesting to find other systems, that show these properties. Finding a general condition, possibly using the topological arguments, which could allow prediction of static spheres and Aschenbach effect, would be nice to explore. Checking the occurrence of Aschenbach effect with a charged test particle for our system and also for dyonic black hole in quasi-topological electromagnetism [134], in addition to studying the effect of charge of the black hole, are possible directions for future work.
There are other interesting avenues for future work. From a theoretical standpoint, photon spheres play an important role in imaging Einstein rings of black holes in AdS [136, 137, 138], in bringing out the conformal symmetries [139, 140, 141] in asymptotically flat spacetime backgrounds[142, 143, 144, 145], as well as leading to interesting implications for warped AdS3 black holes [146], which can be explored in the massive gravity context. In particular, for the investigations of AdS/CMT (condensed matter theory) correspondence [147], the techniques of optical imaging  [136, 137, 148, 138, 149, 150, 151] can be used to find holographic gravitational duals to real materials [152], when there are impurities or defects in the lattice structure, using holographic massive gravity models [73, 120, 128, 129, 130]. There are further intriguing connections between photon spheres and quasinormal modes in the holographic set up [153, 154, 155, 156, 24, 157, 158, 159, 160, 161, 136, 137], as well as, ISCO’s, anamalous dimensions of certain double twist operators in CFT, and the weak gravity conjecture [162, 163, 164, 165, 166, 167, 168, 169], which should also be explored for the class of massive gravity theories considered in this work. We leave these issues for feature work.